OpenStudy (anonymous):

How do you solve -2x - 18 = 6(1 - 2x) + 4x?

OpenStudy (anonymous):

distribute 6 into (1-2x) and solve the equation like you normally would.

OpenStudy (anonymous):

-2x-18=6-12x+4x -2x-18=6-8x -24=-6x

OpenStudy (anonymous):

so x=4

OpenStudy (anonymous):



OpenStudy (anonymous):

2*6=15x+3 12=15x+3 9=15x

OpenStudy (anonymous):

so x= 15/9

OpenStudy (anonymous):

Can you help me on 2 more please?